# Enhancing Diversity in Parallel Agents: A Maximum State Entropy Exploration Story

### v1

@@ B.2 LAB
  • LABDET: A deterministic version where executing an action $a$ under policy $\pi_{\theta}$ always results in the intended movement.
  • LABSTOCH: A stochastic variant where the agent plays the intended action with probability 1p1-p and deviate from it with $p$.
@@
As outlined in the pseudocode of Algorithm  1, in each episode, a batch $|\mathcal{B}|=40$, of $K$ trajectories are experienced by parallel learners. $K$ is fixed during the experiment always to 1, due to the assumption of having access to the minimum size of possible interactions that the agents can have with respect to the environment. To reproduce the policy gradient performance of the single agent, this $K$ should coincide to the number of agent $m$, at which the single is compared to. Indeed, in the 1 with respectively $\{2,6\}$ agents involved, the single agent experiences $K=\{2,6\}$ trajectories, from which the state distribution $d_{n}$ is derived. Over the $10k$ episodes, the parallel agents interact with the environment with finite trajectories of lenght $T$ that:
  • In the ROOMDET and ROOMS environments is set to $8$,
  • While in LABDET and LABSTOCH, due to a bigger state space, is set to $10$.
@@
$\theta$ parameters are updated via gradient ascent maximization, with learning rate $\alpha$ set initially to $0.1$ with exponential decay rate $\lambda$ of 0.999. The training is performed over $5$ different seed $\{0,1,2,42,133\}$.
@@ C.1 Performance of Four Parallel Agents
To further validate our approach, we extend the experimental analysis presented in the main body by including results for four parallel agents. Figure 7 illustrates the performance of this configuration in terms of normalized state entropy (top row) and support size (bottom row) across the four environments: ROOMDET, ROOMS, LABDET, and LABSTOCH.
@@ 2 Agents1 Agent $K^{\prime}=2K$4 Agents1 Agent $K^{\prime}=4K$6 Agents1 Agent $K^{\prime}=6K$
(a) State Entropy (ROOMDET).
(b) State Entropy (ROOMS).
(c) State Entropy (LABDET).
(d) State Entropy (LABSTOCH).
(e) Support (ROOMDET).
(f) Support (ROOMS).
(g) Support (LABDET).
(h) Support (LABSTOCH).
@@ C.2 Dataset Analysis
To further understand the benefits of parallel exploration, we analyze the datasets generated by the learned parametric policies. These datasets are constructed analogously to those in the single-agent case and the random policy baseline. Specifically, for each parallel-agent configuration, we sample $K=m$ trajectories to ensure a consistent comparison across settings. The mean and standard deviation of the state entropy are computed over these trajectories and averaged across the five random seeds used during training. Let’s look deeply at the dataset created from the paralell agents, when interacting with the more complex LAB environment. Figures 8 present heatmaps visualizing the state occupancy distributions for the LAB environments, respectively the deterministic and stochastic case. These heatmaps provide a qualitative comparison of the exploration patterns achieved by single agents, parallel agents, in the two implemented versions.
(a) Parallel Agents (LABDET).
(b) Parallel Agents (LABSTOCH).
(c) Single Agent (LABDET).
(d) Single Agent (LABSTOCH).
Figure 8: Dataset extrapolation based on the LAB environments.
@@ C.4 Diversity Analysis
Figure 10 provides a visual comparison of the learned policies for parallel agents versus single agents in both ROOMDET and LABDET environments. In the single-agent setting (Figures 10a and 10b), the agent tends to develop policies with higher stochasticity in its action choices. This is a consequence of the need to explore a larger portion of the state space alone, necessitating a more randomized approach to discover high-entropy trajectories. In contrast, the parallel agent setting (Figures 10c and 10d) reveals that the centralized update mechanism, based on the aggregate state distribution $d_{n}$, encourages individual agents to specialize in more deterministic policies.
(a) Single Agent (ROOMDET).
(b) Single Agent (LABDET).
(c) Parallel Agents (ROOMDET).
(d) Parallel Agents (LABDET).

### v2

@@ B.2 Maze
  • Maze-det: A deterministic version where executing an action $a$ under policy $\pi_{\theta}$ always results in the intended movement.
  • Maze-stoc: A stochastic variant where the agent plays the intended action with probability 1p1-p and deviate from it with $p$.
@@
As outlined in the pseudocode of Algorithm  5, in each episode, a batch $|\mathcal{B}|=40$, of $K$ trajectories are experienced by parallel learners. $K$ is fixed during the experiment always to 1, due to the assumption of having access to the minimum size of possible interactions that the agents can have with respect to the environment. To reproduce the policy gradient performance of the single agent, this $K$ should coincide to the number of agent $m$, at which the single is compared to. Indeed, in the 1 with respectively $\{2,6\}$ agents involved, the single agent experiences $K=\{2,6\}$ trajectories, from which the state distribution $d_{n}$ is derived. Over the $10k$ episodes, the parallel agents interact with the environment with finite trajectories of lenght $T$ that in the Room-det and Room-stoc environments is set to $8$, while in Maze-det and Maze-stoc, due to a bigger state space, is set to $10$.
@@
$\theta$ parameters are updated via gradient ascent maximization, with learning rate $\alpha$ set initially to $0.1$ with exponential decay rate $\lambda$ of 0.999. The training is performed over $5$ different seeds $\{0,1,2,42,133\}$.
@@ Appendix C Further Experimental Details and Analysis
The code to reproduce the experiments is made available at the following link.
@@ C.1 Performance of Four Parallel Agents
To further validate our approach, we extend the experimental analysis presented in the main body by including results for four parallel agents. Figure 7 illustrates the performance of this configuration in terms of normalized state entropy (top row) and support size (bottom row) across the four environments: Room-det, Room-stoc, Maze-det, and Maze-stoc.
@@ 2 Agents1 Agent $K^{\prime}=2K$4 Agents1 Agent $K^{\prime}=4K$6 Agents1 Agent $K^{\prime}=6K$
(a) State Entropy (Room-det).
(b) State Entropy (Room-stoc).
(c) State Entropy (Maze-det).
(d) State Entropy (Maze-stoc).
(e) Support (Room-det).
(f) Support (Room-stoc).
(g) Support (Maze-det).
(h) Support (Maze-stoc).
@@ C.2 Dataset Analysis
To further understand the benefits of parallel exploration, we analyze the datasets generated by the learned parametric policies. These datasets are constructed analogously to those in the single-agent case and the random policy baseline. Specifically, for each parallel-agent configuration, we sample $K=m$ trajectories to ensure a consistent comparison across settings. The mean and standard deviation of the state entropy are computed over these trajectories and averaged across the five random seeds used during training. Let’s look deeply at the dataset created from the paralell agents, when interacting with the more complex Maze environment. Figures 8 present heatmaps visualizing the state occupancy distributions for the Maze environments, respectively the deterministic and stochastic case. These heatmaps provide a qualitative comparison of the exploration patterns achieved by single agents, parallel agents, in the two implemented versions.
(a) Parallel Agents (Maze-det).
(b) Parallel Agents (Maze-stoc).
(c) Single Agent (Maze-det).
(d) Single Agent (Maze-stoc).
Figure 8: Dataset extrapolation based on the Maze environments.
@@ C.4 Diversity Analysis
Figure 10 provides a visual comparison of the learned policies for parallel agents versus single agents in both Room-det and Maze-det environments. In the single-agent setting (Figures 10a and 10b), the agent tends to develop policies with higher stochasticity in its action choices. This is a consequence of the need to explore a larger portion of the state space alone, necessitating a more randomized approach to discover high-entropy trajectories. In contrast, the parallel agent setting (Figures 10c and 10d) reveals that the centralized update mechanism, based on the aggregate state distribution $d_{n}$, encourages individual agents to specialize in more deterministic policies.
(a) Single Agent (Room-det).
(b) Single Agent (Maze-det).
(c) Parallel Agents (Room-det).
(d) Parallel Agents (Maze-det).